OpenStudy (anonymous):

Solve the following absolute value inequality: |3x - 1| - |x + 1| > 0 I don't understand how to do these forms of equations when there are two absolute value terms. Would anyone mind giving me a hand here?

OpenStudy (anonymous):

\[|3x -1|>|x-1|\]

OpenStudy (bahrom7893):

it's 3x+1>x+1 or -(3x+1)<-(x+1)

OpenStudy (bahrom7893):

isn't it? sorry haha i haven't done abs val inequalities in a while.. i might be wrong

OpenStudy (anonymous):

I am not sure as to how you are guys are able to get rid of the absolute bars

OpenStudy (anonymous):

Ishaan94, how do you know that?

OpenStudy (anonymous):

Ok so you have to basically work the cases for when they are that many units away from the positive side, and that many units away from the negative side.

OpenStudy (anonymous):

And since there are two absolute value arguments here, this means that you have 4 combos?

OpenStudy (anonymous):

actually its 3x+1>x+1 x>0 and 3x+1<-(x+1) 4x<-2 x<-1/2

OpenStudy (anonymous):

So what i said from before is that true? You have to go back to the equation and verify that LS > RS?

OpenStudy (anonymous):

how can that b? x=5 works (3*5-1)-(5-1)>0 15-1-4>0 10>0

OpenStudy (zarkon):

x<0 or x>1

OpenStudy (zarkon):

the main thing here, I believe, is that there are not 4 cases

OpenStudy (anonymous):

wait a moment zarkon has that x<0 and x>1 so 1/2 shouldnt work 3(1/2)+1=5/2 1/2-1=-1/2 so 5/2-1/2>0 4/2>0 2>0 so zarkon is wrong as well

OpenStudy (zarkon):

try that again

OpenStudy (anonymous):

my bad messedd it up u r right

OpenStudy (zarkon):

I am correct

OpenStudy (zarkon):

if (3x-1)>0 then 3x>1 and x>1/3 then x+1>1/3+1>0 also if x+1<0 then x<-1 then 3x-1<0 that kills two of the four cases

OpenStudy (zarkon):

you can't have 3x-1>0 and x+1<0 you can't have x+1<0 and 3x-1>0

OpenStudy (anonymous):

well zarkon is right sorry i did it all wrong ..it's kinda weird i did that way too wrong sorry : (

OpenStudy (anonymous):

yes he is right 4 sure.